# A Survey of Graph Retrieval-Augmented Generation for Customized Large Language Models

### v1

@@ V-B3 GNN-based Retriever
Graph neural networks (GNNs) have become a mainstream tool for graph modeling and mining, achieving significant results in tasks such as node classification [137, 138, 139, 140] and link prediction [141, 142, 143]. The message-passing mechanism in GNNs operates by iteratively aggregating information from a node’s neighbors to update its own representation, enabling the network to learn from the structure and features of the graph [144, 137]. GNN-based knowledge retriever primarily utilizes graph neural networks to encode nodes within the constructed graph bases [31, 62, 57, 58, 57] . The localizing and positioning of knowledge mainly rely on the encoding similarity of the node representations that contain both sentiment meanings and structural relationship understandings. GNN-based retrievers require the training of a GNN encoder. Additionally, due to the lack of explicitly labeled data, the focus of training is on designing an appropriate loss function that enables the GNN to learn to accurately locate the target knowledge through representation encoding.
@@ V-C Retrieval Enhancement Strategies
Based on the retrieval technologies, some strategies are designed to enhance the retrieved resources, aiming to improve the relevance of target queries. We will introduce insights of these strategies in this subsection.
@@ V-C1 Multi-round Retrieval
Beyond static retrieval techniques, some GraphRAG systems incorporate contextual information and user feedback to refine and adapt the retrieval process dynamically, which we call multi-round retrieval. These approaches aim to gradually align the retrieval mechanism more closely with specific use cases and evolving user intents [62, 63, 47].
@@ V-C3 Hybrid Retrieval
Combining different forms of data for retrieval can integrate their strengths and complement their weaknesses. In GraphRAG systems, there are also some recent works that adopt both graphs and other data architectures to collaboratively finish the RAG process. Specifically in recent years, to enhance the scope of the retrieval candidates, Graph+Vector RAG [119, 153] combine the knowledge graph and the vector database as the retrieval sources and integrate the retrieved results for the answer generation. For example, ToG-2 [23] performs joint retrieval on a knowledge graph and text documents. It begins with entities identified from the query and iteratively retrieves relevant knowledge. Due to the static nature of knowledge graphs, to further ensure the timeliness and accuracy of the retrieved information, some efforts have begun to combine knowledge graphs with online web resources for joint retrieval, which can be summarized as Graph+Online Web Resource RAG [154, 155, 156, 157, 158, 159]. Further, K2 [160] explores the LLM tuning with multiple retrieval sources, such as the textual corpus, self-instructed LLM knowledge, and online web resources. In future works, hybrid RAG with multiple retrieval resources simultaneously can be a potential research direction for complex LLM system building under a wide range of knowledge.
TABLE III: Representative fine-tuning knowledge integration techniques used in GraphRAG systems.
@@
V-D Discussions
The true magic of graph-based knowledge resources lies in their supernatural ability to connect data while being flexible and scalable during retrieval. Graphs can seamlessly integrate various data sources, including structured and unstructured data. This fusion creates a unified view of information, often revealing hidden connections and patterns. Advanced graph knowledge bases incorporate ontologies and semantic schemas, providing formal definitions for concepts and their interrelationships within a domain. While traditional retrieval methods are like tourists asking random pedestrians for directions, GraphRAG is the mayor who knows every nook and cranny of the city. It doesn’t just look at individual buildings (or pieces of information); it sees the entire cityscape, considering the intricate network of streets, neighborhoods, and communities. This bird’s-eye view allows GraphRAG to navigate these streets with the efficiency of a seasoned cab driver who knows all the shortcuts. Here is a vivid example: Let’s consider GraphRAG to be a sophisticated urban planning and information retrieval system for the knowledge base aka city. This system operates in two main phases: planning and retrieval.
In the planning phase, the city planners (LLMs) first divide the urban landscape (source documents) into manageable districts (concepts). They then survey each district, identifying key landmarks and connections (element instances) - think of this as mapping out important buildings, parks, and the roads connecting them. These individual elements are then summarized into concise descriptions (element summaries), like creating brief profiles of each neighborhood landmark. Next, using advanced urban analysis tools (community detection algorithms like Leiden), the planners group these elements into natural communities, much like identifying distinct neighborhoods or boroughs within the city. Finally, they craft detailed ”neighborhood reports” (community summaries) for each of these communities, providing a rich, multi-layered guidebook to the entire city. This indexing phase creates a comprehensive, hierarchical understanding of the urban landscape, from individual buildings to entire districts.
After the planning, we obtain a high-quality graph knowledge base. Then, when it comes to retrieval, GraphRAG shines in handling citywide inquiries. Instead of dispatching surveyors to random locations, it consults the pre-prepared neighborhood reports. Each community contributes its local perspective to the broader question, like local town halls holding simultaneous meetings (community summaries to community answers). These local insights are then synthesized into a comprehensive city report (global answer), much like an urban planner combining feedback from various boroughs to understand citywide trends. This approach proved particularly effective for ”global” questions about the entire urban landscape, outperforming traditional methods in both the breadth and diversity of insights offered.
@@ VI Knowledge Integration
The knowledge retrieval phase is crafted to gather pertinent documents from various external sources in alignment with a specified query. Following this, the integration phase focuses on seamlessly synthesizing documents obtained from knowledge retrieval into a cohesive prompt, simultaneously setting appropriate training goals for the purpose of optimization. Recognizing that LLMs form a sturdy and foundational framework, the exploration during the integration phase predominantly steers clear of modifying the internal structure of individual layers or necessitating a complete reinitialization of LLMs for training. In this section, we delve into the comprehensive pipeline, relevant technologies, and strategies for enhancing knowledge integration.
@@ VI-B Integration Techniques
TABLE IV: Representative in-context learning knowledge integration techniques used in GraphRAG systems.
@@ VI-B2 In-context Learning Techniques
Graph-enhanced Chain-of-Thought. Since Chain-of-Thought (CoT) has demonstrated effective performance improvement and scalability [180], CoT-based prompting techniques on knowledge integration also elicit strong potential in complex multi-hop reasoning. The Graph-enhanced CoT, such as Reasoning on Graphs [42], emphasizes chain reasoning based on the interaction between graphs and LLMs. Think-on-Graph [22, 23] proposes a reasoning framework based on the KG-LLMs interaction; LLMs are required to retrieve the relevant sub-graph of the proposed keyword first, then deduce the final result through chain reasoning. Similarly, in Graph CoT [63], given a question, LLM reasons the key component in the external graph that is required for the final answer, then uses a multi-round graph execution and interaction to finish the chain reasoning that concludes the final answer. Chain-of-Knowledge [75], firstly constructs exemplars that answer the given question based on evidence triples retrieved from the knowledge base, then utilizes faithfulness and factuality verification to check the reliability of the chain prediction. Program-of-Thought (PoT) [181, 182] is also a well-designed prompting technique that extends CoT prompting into coding scenarios. Following this, the KnowledGPT [77] framework uses PoT to generate the research language of knowledge bases to retrieve the answer step by step. The CoT-based methods achieve promising progress in knowledge integration for GraphRAG, significantly improving the reasoning capability of complex questions for LLMs. However, error accumulation is still a severe problem in this ICL paradigm [183, 184]. To address this, research is being conducted to diversify the reasoning paths that LLMs can take to reach the same conclusion [185, 186], which can help mitigate the effects of error propagation. Additionally, the integration of visual in-context learning for Large Vision-Language Models is being explored to enhance cross-modal interactions and representation alignment [187, 188], which may provide further insights into managing error accumulation in reasoning tasks.
@@
Collaborative Knowledge Graph Refinement. Instead of a well-designed prompting technique utilized to generate a better prompt or elicit LLMs for better reasoning, refining the LLM’s original response based on the factual knowledge in knowledge graphs is also an effective method to prevent LLMs from hallucination scenarios. The timeliness and accuracy of knowledge graphs are crucial, as they greatly influence the quality of augmented generation [63, 189, 190]. Maintaining and refining the content in knowledge graphs through feedback in LLM generation is an important strategy in GraphRAG. Representatively, KG-based Retrofitting [43] is a framework that combines LLMs with KGs to mitigate hallucination during the reasoning process. It retrofits the initial draft responses of LLMs based on the factual knowledge stored in KGs with an autonomous knowledge verifying and refining procedure. Similarly, KELP [44] refines a trainable encoder for path selection based on the LLM’s response. The Explore-then-Determine framework [79] uses the integration of LLMs and KGs to determine the final answer. CogMG [80] has designed a framework for collaborative augmentation between LLMs and KGs, leveraging KGs to augment LLMs in generations, explicitly targeting incomplete knowledge. LLMs are then required to identify and decompose required knowledge triples that are not present in the KG, enriching them and aligning updates with real-world demands. The collaboration between KGs and LLMs aligns factual knowledge to ensure the accuracy of generation and the up-to-date quality of KGs. However, ensuring the correctness of the intermediate process is challenging, and the accuracy of the refining process should also be evaluated with tailored metrics [191, 192, 193].
@@ VI-C1 Training with other domain-specific models
Training with domain-specific models is a critical approach to enhance the capabilities of LLMs in handling multi-modal information representations and to improve their performance on specific tasks. (i) Enhancing Multi-modal Information Processing: Integrating multi-modal information representations, such as image representations learned by vision-language models [194], into LLMs, is a significant advancement. These representations not only reduce input lengths but also encapsulate rich information from specific inputs. However, they also pose challenges for LLMs, as they may be treated as out-of-vocabulary tokens, hindering the full exploitation of the original multi-modal information. Furthermore, recent efforts demonstrate the large number of additional tokens especially in the extremely sparse case where each additional token is only paired with a few training samples can cause LLMs to perform even worse over the small LMs [195]. To address this, recent efforts have focused on fine-tuning LLMs with domain-specific models to increase their generalization capabilities across different domains [196]. (ii) Fine-tuning LLM with Domain-specific Task: Fine-tuning an LLM involves using a pre-trained model and refining its weights by training it on a small set of annotated data with a slow learning rate. This principle allows the language model to adopt new knowledge from the data while retaining its initial learnings [197]. Transfer learning, a technique that allows a pre-trained model to apply its knowledge to a new task, is particularly useful when sufficient datasets for fine-tuning are not available [198]. For instance, MedPaLM [199] is a domain-specific model built upon PaLM, which demonstrated exceptional performance in complex tasks by using prompting strategies with annotated medical questions and answers.

### v2

@@ V-B3 GNN-based Retriever
Graph neural networks (GNNs) have become a mainstream tool for graph modeling and mining, achieving significant results in tasks such as node classification [152, 153, 154, 155] and link prediction [156, 157, 158]. The message-passing mechanism in GNNs operates by iteratively aggregating information from a node’s neighbors to update its own representation, enabling the network to learn from the structure and features of the graph [152]. GNN-based knowledge retriever primarily utilizes graph neural networks to encode nodes within the constructed graph bases [44, 92, 85, 85] . The localizing and positioning of knowledge mainly rely on the encoding similarity of the node representations that contain both sentiment meanings and structural relationship understandings. GNN-based retrievers require the training of a GNN encoder. Additionally, due to the lack of explicitly labeled data, the focus of training is on designing an appropriate loss function that enables the GNN to learn to accurately locate the target knowledge through representation encoding.
@@ V-C Retrieval Enhancement Strategies
Based on the retrieval technologies, some strategies are designed to enhance the retrieved resources, aiming to improve the relevance of target queries. We will introduce insights into these strategies in this subsection.
@@ V-C1 Multi-round Retrieval
Beyond static retrieval techniques, some GraphRAG systems incorporate contextual information and user feedback to refine and adapt the retrieval process dynamically, which we call multi-round retrieval. These approaches aim to gradually align the retriever more closely with specific use cases and evolving user intents [92, 93, 73].
@@ V-C3 Hybrid Retrieval
Combining different forms of data for retrieval can integrate their strengths and complement their weaknesses. In GraphRAG systems, recent studies have adopted the use of both graphs and other data formats to collaboratively enhance the RAG process. Specifically, to enhance the scope of the retrieval candidates, Graph+Vector RAG [135, 167] combine the knowledge graph and the vector database as the retrieval sources and integrate the retrieved results for answer generation. For example, ToG-2 [22] performs joint retrieval on a knowledge graph and text documents. It begins with entities identified from the query and iteratively retrieves relevant knowledge. Due to the static nature of knowledge graphs, to further ensure the timeliness and accuracy of the retrieved information, some efforts have begun to combine knowledge graphs with online web resources for joint retrieval, which can be summarized as Graph+Online Web Resource RAG [168, 169, 170, 171, 172, 173]. Further, K2 [174] explores the LLM tuning with multiple retrieval sources, such as the textual corpus, self-instructed LLM knowledge, and online web resources. In future works, hybrid RAG with multiple retrieval resources simultaneously can be a potential research direction for complex LLM system building under a wide range of knowledge.
TABLE II: Representative fine-tuning knowledge integration techniques used in GraphRAG systems.
@@ VI Knowledge Integration
The knowledge retrieval phase is crafted to gather pertinent documents from various external sources in alignment with a specified query. Following this, the integration phase focuses on seamlessly synthesizing documents obtained from knowledge retrieval into a cohesive prompt, simultaneously setting appropriate training goals for the purpose of optimization. Recognizing that LLMs form a sturdy and foundational framework, the exploration during the integration phase predominantly steers clear of modifying the internal structure of individual layers or necessitating a complete reinitialization of LLMs for training. In this section, we delve into the comprehensive pipeline, relevant technologies, and strategies for enhancing knowledge integration in different GraphRAG systems.
@@ VI-B Integration Techniques
VI-B1 Fine-tuning Techniques
TABLE III: Representative in-context learning knowledge integration techniques used in GraphRAG systems.
@@ VI-B2 In-context Learning Techniques
Graph-enhanced Chain-of-Thought. Since Chain-of-Thought (CoT) has demonstrated effective performance improvement and scalability [194], CoT-based prompting techniques on knowledge integration also elicit strong potential in complex multi-hop reasoning. The Graph-enhanced CoT, such as Reasoning on Graphs [65], emphasizes chain reasoning based on the interaction between graphs and LLMs. Think-on-Graph [21, 22] proposes a reasoning framework based on the KG-LLMs interaction; LLMs are required to retrieve the relevant sub-graph of the proposed keyword first, then deduce the final result through chain reasoning. Similarly, in Graph CoT [93], given a question, LLM reasons the key component in the external graph that is required for the final answer, then uses a multi-round graph execution and interaction to finish the chain reasoning that concludes the final answer. Chain-of-Knowledge [107], firstly constructs exemplars that answer the given question based on evidence triples retrieved from the knowledge base, then utilizes faithfulness and factuality verification to check the reliability of the chain prediction. Program-of-Thought (PoT) [195, 196] is also a well-designed prompting technique that extends CoT prompting into coding scenarios. Following this, the KnowledGPT [109] framework uses PoT to generate the research language of knowledge bases to retrieve the answer step by step. The CoT-based methods achieve promising progress in knowledge integration for GraphRAG, significantly improving the reasoning capability on complex questions. However, error accumulation is still a severe problem [197, 198]. To address this, research has been conducted to diversify the reasoning paths that LLMs can take to reach the same conclusion [199, 200], which can help mitigate the effects of error propagation. Besides, the integration of visual in-context learning for Large Vision-Language Models is being explored to enhance cross-modal interactions and alignment [201, 202], which may provide further insights into managing error accumulation during reasoning.
@@
Collaborative Knowledge Graph Refinement. Instead of a well-designed prompting technique utilized to generate a better prompt or elicit LLMs for better reasoning, refining the LLM’s original response based on the factual knowledge in knowledge graphs is also an effective method to prevent LLMs from hallucination scenarios. The timeliness and accuracy of knowledge graphs are crucial, as they greatly influence the quality of augmented generation [93]. Maintaining and refining the content in knowledge graphs through feedback in LLM generation is an important strategy in GraphRAG. Representatively, KG-based Retrofitting [66] is a framework that combines LLMs with KGs to mitigate hallucination during the reasoning process. It retrofits the initial draft responses of LLMs based on the factual knowledge stored in KGs with an autonomous knowledge verifying and refining procedure. Similarly, KELP [67] refines a trainable encoder for path selection based on the LLM’s response. The Explore-then-Determine framework [111] uses the integration of LLMs and KGs to determine the final answer. CogMG [112] has designed a framework for collaborative augmentation between LLMs and KGs, leveraging KGs to augment LLMs in generations, explicitly targeting incomplete knowledge. LLMs are then required to identify and decompose required knowledge triples that are not present in the KG, enriching them and aligning updates with real-world demands. The collaboration between KGs and LLMs aligns factual knowledge to ensure the accuracy of generation and the up-to-date quality of KGs. However, ensuring the correctness of the intermediate process is challenging, and the accuracy of the refining process should also be evaluated with tailored metrics [203].
@@ VI-C1 Training with other domain-specific models
Training with domain-specific models is a critical approach to enhance the capabilities of LLMs in handling multi-modal information representations and to improve their performance on specific tasks. (i) Enhancing Multi-modal Information Processing: Integrating multi-modal information representations, such as image representations learned by vision-language models [204], into LLMs, is a significant advancement. These representations not only reduce input lengths but also encapsulate rich information from specific inputs. However, they also pose challenges for LLMs, as they may be treated as out-of-vocabulary tokens, hindering the full exploitation of the original multi-modal information. Furthermore, recent efforts demonstrate the large number of additional tokens, especially in the extremely sparse case where each additional token is only paired with a few training samples, can cause LLMs to perform even worse over the small LMs. To address this, recent efforts have focused on fine-tuning LLMs with domain-specific models to increase their generalization capabilities across different domains [205]. (ii) Fine-tuning LLM with Domain-specific Tasks: Fine-tuning an LLM involves using a pre-trained model and refining its weights by training it on a small set of annotated data with a slow learning rate. This principle allows the language model to adopt new knowledge from the data while retaining its initial learnings [206]. Transfer learning, a technique that allows a pre-trained model to apply its knowledge to a new task, is particularly useful when sufficient datasets for fine-tuning are not available [207]. For instance, MedPaLM [208] is a domain-specific model built upon PaLM, which demonstrated exceptional performance in complex tasks by using prompting with annotated questions and answers.